OpenStudy (1645323):

Find the focus of the parabola y = ⅛(x - 3)2 + 1. Select one: a. (-3, -1) b. (5, 1) c. (-1, 3) d. (3, 3)

OpenStudy (anonymous):

(x-h)^2 = 4p(y-k), where p is the distance from the vertex to the focus

OpenStudy (anonymous):

find the vertex, and find p, then you'll have the focus

OpenStudy (1645323):

SO is it C? @sourwing

OpenStudy (jdoe0001):

what did you get for the vertex?

OpenStudy (1645323):

I just got a line not a parabola @jdoe0001

OpenStudy (jdoe0001):

hmm that looks like a parabola -> \(\bf y=\cfrac{1}{8}(x-3)^2+1\)

OpenStudy (jdoe0001):

can you tell the vertex from it?

OpenStudy (jdoe0001):

ttp://www.mathwarehouse.com/geometry/parabola/images/standard-vertex-forms.png

OpenStudy (jdoe0001):

shoot... missing letter =) http://www.mathwarehouse.com/geometry/parabola/images/standard-vertex-forms.png

OpenStudy (1645323):

nevermind I got the parabola

OpenStudy (1645323):

3,1

OpenStudy (1645323):

is the vertex @jdoe0001

OpenStudy (jdoe0001):

yeap and "p" as @sourwing already pointed out, is the distance from the vertex to the focus the parabola is "x" based, so is vertical so we know the parabola has a POSITIVE leading term coefficient that means the parabola is going UP so from the vertex of (3,1) it goes up, meaning the "y" value moves how far up? well, "p" distance \(\bf (x-{\color{red}{ h}})^2=4{\color{green}{ p}}(y-{\color{blue}{ k}})\qquad vertex\quad ({\color{red}{ h}},{\color{blue}{ k}}) \\ \quad \\ \quad \\ y=\cfrac{1}{8}(x-3)^2+1\implies y-1=\cfrac{1}{8}(x-3)^2 \\ \quad \\ \implies 8(y-1)=(x-3)^2\) can you tell "p" from there?

OpenStudy (1645323):

.5

OpenStudy (jdoe0001):

well \(\bf (x-{\color{red}{ h}})^2=4{\color{green}{ p}}(y-{\color{blue}{ k}})\qquad vertex\quad ({\color{red}{ h}},{\color{blue}{ k}}) \\ \quad \\ \quad \\ \implies 8(y-1)=(x-3)^2\implies 4{\color{green}{ p}}=8\implies p=?\)

OpenStudy (1645323):

2 nevermind

OpenStudy (jdoe0001):

yeap :), so the focus is at (3, 1 + 2 )

OpenStudy (1645323):

so 3,3?

OpenStudy (jdoe0001):

yeap